# Fin-RATE: A Real-world Financial Analytics and Tracking Evaluation Benchmark for LLMs on SEC Filings

### v1

@@ Abstract.
With increasing deployment of Large Language Models (LLMs) in the finance domain, LLMs are increasingly expected to parse complex regulatory disclosures. However, existing benchmarks often focus on isolated details, failing to reflect the complexity of professional analysis that requires synthesizing information across multiple documents, reporting periods, and corporate entities. They do not distinguish whether errors stem from retrieval failures, generation flaws, finance-specific reasoning mistakes, or misunderstanding of the query or context. This makes it difficult to pinpoint performance bottlenecks. To bridge these gaps, we introduce Fin-RATE, a benchmark built on U.S. Securities and Exchange Commission (SEC) filings and mirror financial analyst workflows through three pathways: detail-oriented reasoning within individual disclosures, cross-entity comparison under shared topics, and longitudinal tracking of the same firm across reporting periods. We benchmark 17 leading LLMs, spanning open-source, closed-source, and finance-specialized models, under both ground-truth context and retrieval-augmented settings. Results show substantial performance degradation, with accuracy dropping by 18.60% and 14.35% as tasks shift from single-document reasoning to longitudinal and cross-entity analysis. This is driven by rising comparison hallucinations, time and entity mismatches, and mirrored by declines in reasoning and factuality—limitations that prior benchmarks have yet to formally categorize or quantify.
@@ 1. Introduction
In recent years, the application of general-purpose large language models (LLMs) in the financial domain has rapidly expanded, with an increasing number of financial institutions exploring their use in high-value tasks such as financial report analysis, risk monitoring, and investment research assistance. According to the McKinsey 2025 report, several finance teams have already leveraged LLMs to improve the accuracy of budgeting and the timeliness of financial reporting. LLMs are viewed as promising analytical assistants, expected to undertake parts of professional analysis workflows, particularly in parsing structurally complex and semantically dense regulatory disclosures such as SEC filings. However, effective understanding of such documents goes far beyond fact-level extraction or isolated question answering. Instead, it involves cross-document, cross-temporal, and cross-entity information integration and structured reasoning, imposing significantly higher demands on a model’s consistency, depth of reasoning, information aggregation, and robustness in expression.
Table 1. The criteria columns of the table include: (1) cross-company comparison, (2) cross-year tracking, (3) multi-document reasoning, (4) analysis-oriented question formulation, (5) fine-grained evaluation beyond answer-level correctness, and (6) data source coverage. Fin-RATE incorporates a wide range of SEC filings beyond 10-K, 10-Q, and 8-K (see Appendix D for details).
@@
While prior studies have shown that LLMs exhibit promising capabilities in financial knowledge QA and structured numerical reasoning (Guo et al., 2025b; Tang et al., 2025), multiple benchmarks grounded in real-world financial filings (Islam et al., 2023; Reddy et al., 2024; Choi et al., 2025; Lai et al., 2025; Zhang et al., 2025; Ji et al., 2025) have highlighted persistent limitations in several core dimensions: (1) difficulty in integrating dispersed information from long documents and establishing reasoning paths across paragraphs and time periods; (2) challenges in maintaining factual consistency between answers and the original disclosures, especially under complex semantic contexts; and (3) absence of causal and explanatory reasoning to justify comparisons, interpret trends, or verify claims based on financial disclosures.
Although these limitations have been partially identified in prior benchmarks, existing evaluations still suffer from systematic shortcomings in both task design and evaluation methodology. (1) Most benchmarks oversimplify SEC filings as flat, isolated text spans, focusing on information extraction, numerical computation over tabular data, and QA restricted to a single section and reporting year (Reddy et al., 2024; Chen et al., 2021; Islam et al., 2023; Spörer, 2025). This narrow view ignores competitive dynamics between peer firms, the evolution of a single company across years, and the mutually reinforcing relationship between sections. As a result, the depth and breadth of the constructed questions fall far short of real-world analyst workflows, failing to effectively evaluate LLMs for practical financial analysis. (2) Evaluation protocols remain coarse-grained, focusing primarily on final answer correctness. Even benchmarks that incorporate LLM-as-Judge modules(Choi et al., 2025; BenYoash et al., 2025), remain confined to generic axes like faithfulness. They provide little visibility into how and why models degrade on complex financial analysis, struggling to quantify capability loss or reveal concrete failure modes.
@@
To close these gaps, we propose Fin-RATE, a benchmark designed to model multi-dimensional alignment across corporate entities, temporal sequences, and heterogeneous disclosure structures as shown in Figure 1. Detail & Reasoning QA (DR-QA) assesses if a model can extract and integrate related information within a financial disclosure section to interpret specific financial impacts or perform calculations. Enterprise Comparison QA (EC-QA) evaluates the model’s ability to align disclosures from different companies and identify salient differences across structured metrics and narrative content, despite variations in reporting style. Longitudinal Tracking QA (LT-QA) examines whether the model can detect trends and shifts across filings from the same company over time, including structural or semantic changes under evolving disclosure frames. To systematically diagnose model failures, we construct a fine-grained error taxonomy derived from extensive manual examination of model outputs on SEC filing tasks. Based on recurring failure patterns, we define thirteen core types across four categories: retrieval errors, generation inconsistencies, financial reasoning failures, and comprehension issues. We summarize our key contributions as follows:
@@
  • We evaluate 17 LLMs under gold-context and retrieval-augmented settings, observing distinct capability profiles. Finance-specific models perform well in numeric reasoning and domain terminology but break down under complex reasoning demands.
@@ 2. Related Work
Early financial QA benchmarks, such as FinQA (Chen et al., 2021), TAT-QA (Zhu et al., 2021), and PACIFIC (Deng et al., 2022), rely on curated tables and text excerpts, treating financial information as isolated knowledge units. Such atomization reduces analysis to symbolic computation, shielding models from the semantic noise and narrative complexity inherent in full SEC filings. With the advancement of long-context language models, subsequent benchmarks including FinanceBench (Islam et al., 2023), FinDER (Choi et al., 2025), and DocFinQA (Reddy et al., 2024) began to incorporate full-length SEC disclosure documents. These works focus on evaluating information retrieval and factual extraction under long-document settings. Despite incorporating richer contexts, these benchmarks still fundamentally model filings as flattened databases, where core tasks are limited to point-wise fact lookup or single-period detail verification. As a result, they do not test whether models can reason about logical consistency, structural alignment, or semantic evolution across disclosures. To extend beyond single-document retrieval, SEC-QA (Lai et al., 2025) introduces cross-company queries through a template-based generation pipeline. While enabling limited cross-entity comparison, the resulting questions are highly rigid in structure and struggle to capture the heterogeneous and narrative-driven nature of disclosures across different companies. Although SECQUE (BenYoash et al., 2025) further incorporates expert-authored analytical questions spanning tasks such as trend analysis, its comparative settings focus on financial metrics and remain restricted to single reporting periods, leaving non-financial dimensions like risk factors largely unexplored. Beyond task construction, existing benchmarks also suffer from a pronounced black-box issue in their evaluation mechanisms. As illustrated in Table 1, prior work relies solely on answer-level accuracy or coarse-grained labels such as Correct, Partial, or Incorrect. Such shallow evaluation protocols obscure the true failure modes of language models in complex financial reasoning scenarios, making it difficult to distinguish whether errors stem from long-document perception degradation, breakdowns in multi-step financial logic, or misalignment across temporal and structural contexts.
@@ 3.1. Data Source and Statistics
As illustrated in Figure 2, Fin-RATE is built entirely from official filings retrieved via the U.S. SEC EDGAR database [^1] https://www.sec.gov/edgar/search-and-access, ensuring authoritative, traceable, and consistent data sourcing. We select a representative set of filing types that together cover the key dimensions required for realistic financial analysis, including periodic reports (10-K, 10-Q), event-driven disclosures (8-K), governance materials (DEF 14A), foreign issuer updates (6-K), equity incentive plans (S-8), securities offerings (424B series), and shareholder communications (SC series). To support reasoning across firms and over time, we further filter companies and filings to obtain a corpus with diverse business models, regulatory exposures, and risk profiles, evaluating model robustness across the heterogeneous reporting styles. The collection spans sectors ranging from relatively stable industries such as utilities and consumer goods to dynamic or newly regulated areas such as cryptocurrency services and biopharmaceutical R&D.
All the filings are converted into Markdown, retaining section headings, tables, and lists. We segment each filing with SEC-standard item boundaries into coherent analytical units to support reasoning and retrieval at multiple levels of granularity. The final corpus comprises 15,311 document chunks derived from 2,472 filings, covering 43 companies across 36 industries from 2020 to 2025, with each chunk containing an average of 2,649 words. Periodic and event-driven reports like 10-K, 8-K, and 10-Q forms constitute the majority of the collection, with their distributions and regulatory functions detailed in Appendix D. Based on this corpus, we construct three QA tasks: Detail Reasoning (DR-QA), Enterprise Comparison (EC-QA), and Longitudinal Tracking (LT-QA). Each task contains 2,500 instances. DR-QA questions are grounded in a single chunk, while EC-QA and LT-QA require aggregating evidence from an average of 5.6 and 3.0 chunks, respectively.
@@ 3.2.1. Detail & Reasoning QA (DR-QA)
To ensure the professionalism and analytical depth of each generated QA pair, we design adaptive prompt templates guided by the official section definitions of various SEC filing types. Prompts for 10-K filings emphasize annual ratio analysis, 10-Q filings highlight quarterly trend dynamics, and 8-K filings direct attention to the quantitative impact of material events. This ensures that each question focuses on the most analytically salient components of the disclosure, effectively reducing the generation of low-value questions and establishing a high-fidelity benchmark for local financial comprehension. See Appendix A for details.
@@ 3.2.2. Enterprise Comparison QA (EC-QA)
To ensure interpretability and regulatory consistency, EC-QA is framed within the same GICS industry category and limited to companies filing 10-K reports in the same fiscal year, ensuring consistency in disclosure frameworks and data granularity. During text selection, we pinpoint semantically aligned sections using structured headings from SEC filings, prioritizing high-comparability modules such as “Item 1A: Risk Factors” and “Item 7: Management Discussion and Analysis.” Besides, cross-sectional combinations are introduced to support multi-dimensional comparative questions across business operations, risk exposures, governance structures, and financial performance. To mitigate the risks of information dilution and hallucination caused by long input contexts, EC-QA construction process adopts a two-stage generation strategy. It first generates structured summaries for each selected company section and then constructs comparison-oriented QA pairs from these aligned summaries. In this way, EC-QA balances analytical depth with input complexity control. See Appendix  B for details.
@@ 3.2.3. Longitudinal Tracking QA (LT-QA)
LT-QA is constructed based on long-form disclosures, primarily annual 10-K reports and DEF 14A proxy statements, and adopts a two-stage, summary-first generation pipeline similar to that used in EC-QA. First, structured summaries are generated for each company–year–section unit, capturing key aspects such as strategic orientation. Subsequently, QA instances are constructed based on these summaries, with each question explicitly linking the same section across multiple years and reflecting temporal comparisons and evolutionary reasoning. See Appendix C for details.
@@ 3.3. Rigorous Quality Control
Fin-RATE includes three quality control stages designed to ensure the reliability, factual grounding, and financial relevance of all QAs. In the automated quality control phase, we adopt a dual-model generation-verification mechanism. First, the generation model (DeepSeek-V3.2) produces multiple candidate QAs under either chunk-level or document-level contexts, based on different prompting strategies. Then, two structurally distinct and high-performing LLMs (DeepSeek-V3.2 and Qwen3-235B) independently assess each candidate QA, focusing on whether the answer is sufficiently supported by the context, whether it has numerical or logical errors, and whether it relies on unsupported inference. Based on task-specific validation rules and a standardized rubric (see Appendix M), each QA is labeled as Correct, Partial, Incorrect, or Failure. Any QA flagged as erroneous by either model is discarded, and only those rated as correct proceed to the human review stage. In the final manual review phase, each QA that passed automated filtering is confirmed one by one to ensure strict adherence to the original disclosure without subjective interpretation. For each QA type, this process is repeated iteratively until 2,500 valid QAs are retained. Full criteria for manual review are detailed in Appendix H. For QAs involving multiple chunks or cross-document contexts, we examine whether each referenced chunk contributes substantively to the reasoning chain. If any referenced chunk is irrelevant, the QA is either rewritten or discarded. For questions involving numerical values, we manually recalculate them to eliminate potential errors.
@@ 4.1.1. Models
  • Finance-Specific Models: To assess the effect of domain adaptation, we evaluate TouchstoneGPT-7B (Wu et al., 2025), Fin-R1 (Liu et al., 2025a), Fino1-14B (Qian et al., 2025), and FinanceConnect-13B (CeADAR, 2023).

### v2

@@ Abstract.
With the increasing deployment of Large Language Models (LLMs) in the finance domain, LLMs are increasingly expected to parse complex regulatory disclosures. However, existing benchmarks often focus on isolated details, failing to reflect the complexity of professional analysis that requires synthesizing information across multiple documents, reporting periods, and corporate entities. Furthermore, these benchmarks do not disentangle whether errors arise from retrieval failures, generation inaccuracies, domain-specific reasoning mistakes, or misinterpretation of the query or context, making it difficult to precisely diagnose performance bottlenecks. To bridge these gaps, we introduce Fin-RATE, a benchmark built on U.S. Securities and Exchange Commission (SEC) filings and mirroring financial analyst workflows through three pathways: detail-oriented reasoning within individual disclosures, cross-entity comparison under shared topics, and longitudinal tracking of the same firm across reporting periods. We benchmark 17 leading LLMs, spanning open-source, closed-source, and finance-specialized models, under both ground-truth context and retrieval-augmented settings. Results show substantial performance degradation, with accuracy dropping by 18.60% and 14.35% as tasks shift from single-document reasoning to longitudinal and cross-entity analysis. This degradation is driven by increased comparison hallucinations, temporal and entity mismatches, and is further reflected in declines in reasoning quality and factual consistency—limitations that existing benchmarks have yet to formally categorize or quantify.
@@ 1. Introduction
In recent years, the application of general-purpose large language models (LLMs) in the financial domain has rapidly expanded, with an increasing number of financial institutions exploring their use in high-value tasks such as financial report analysis, risk monitoring, and investment research assistance. According to the McKinsey 2025 report (mckinsey2025_finance_ai), several finance teams have already leveraged LLMs to improve the accuracy of budgeting and the timeliness of financial reporting. LLMs are viewed as promising analytical assistants, expected to undertake parts of professional analysis workflows, particularly in parsing structurally complex and semantically dense regulatory disclosures such as SEC filings. However, effective understanding of such documents goes far beyond fact-level extraction or isolated question answering. Instead, it involves cross-document, cross-temporal, and cross-entity information integration and structured reasoning, imposing significantly higher demands on a model’s consistency, depth of reasoning, information aggregation, and robustness in expression.
Table 1. The criteria columns of the table include: (1) cross-company comparison, (2) cross-year tracking, (3) multi-document reasoning, (4) analysis-oriented question formulation, (5) fine-grained evaluation beyond answer-level correctness, and (6) data source coverage. Fin-RATE incorporates a wide range of SEC filings beyond 10-K, 10-Q, and 8-K (see Appendix A for details).
@@
While prior studies have shown that LLMs exhibit promising capabilities in financial knowledge question answering (QA) and structured numerical reasoning (Guo et al., 2025b; Tang et al., 2025), multiple benchmarks grounded in real-world financial filings (Islam et al., 2023; Reddy et al., 2024; Choi et al., 2025; Lai et al., 2025; Zhang et al., 2025; Ji et al., 2025) have highlighted persistent limitations in several core dimensions: (1) difficulty in integrating dispersed information from long documents and establishing reasoning paths across paragraphs and time periods; (2) challenges in maintaining factual consistency between answers and the original disclosures, especially under complex semantic contexts; and (3) absence of explanatory reasoning to justify comparisons, interpret trends, or verify claims based on financial disclosures.
Although these limitations have been partially identified in prior benchmarks, existing evaluations still suffer from systematic shortcomings in both task design and evaluation methodology. (1) Most benchmarks oversimplify SEC filings as flat, isolated text spans, focusing on information extraction, numerical computation over tabular data, and QA restricted to a single section and reporting year (Reddy et al., 2024; Chen et al., 2021; Islam et al., 2023; Spörer, 2025). This narrow view ignores competitive dynamics between peer firms, the evolution of a single company across years, and the mutually reinforcing relationship between sections. As a result, the depth and breadth of the constructed questions fall far short of real-world analyst workflows, failing to effectively evaluate LLMs for practical financial analysis. (2) Evaluation protocols remain coarse-grained, focusing primarily on final answer correctness. Even benchmarks that incorporate LLM-as-Judge modules (Choi et al., 2025; BenYoash et al., 2025) remain confined to generic axes like faithfulness. They provide little visibility into how and why models degrade on complex financial analysis, struggling to quantify capability loss or reveal concrete failure modes.
@@
To close these gaps, we propose Fin-RATE, a benchmark designed to model multi-dimensional alignment across corporate entities, temporal sequences, and heterogeneous disclosure structures as shown in Figure 1. Detail & Reasoning QA (DR-QA) assesses if a model can extract and integrate related information within a financial disclosure section to interpret specific financial impacts or perform calculations. Enterprise Comparison QA (EC-QA) evaluates the model’s ability to align disclosures from different companies and identify salient differences across structured metrics and narrative content, despite variations in reporting style. Longitudinal Tracking QA (LT-QA) examines whether the model can detect trends and shifts across filings from the same company over time, including structural or semantic changes under evolving disclosure frames. Furthermore, to systematically diagnose model failures, we construct a fine-grained error taxonomy derived from extensive manual examination of model outputs on SEC filing tasks. Based on recurring failure patterns, we define thirteen core types across four categories: retrieval errors, generation inconsistencies, financial reasoning failures, and comprehension issues. We further observe persistent entity- and year-level retrieval mismatches in EC-QA and LT-QA, motivating the hierarchical retrieval strategy introduced later. We summarize our key contributions below:
@@
  • We evaluate 17 LLMs under ground-truth context and retrieval-augmented settings, observing distinct capability profiles. Finance-specific models perform well in numeric reasoning and domain terminology but break down under complex reasoning demands.
@@ 2. Related Work
Early financial QA benchmarks, such as FinQA (Chen et al., 2021), TAT-QA (Zhu et al., 2021), and PACIFIC (Deng et al., 2022), rely on curated tables and text excerpts, treating financial information as isolated knowledge units. Such atomization reduces analysis to symbolic computation, shielding models from the semantic noise and narrative complexity inherent in full SEC filings. With the advancement of long-context language models, subsequent benchmarks, including FinanceBench (Islam et al., 2023), FinDER (Choi et al., 2025), and DocFinQA (Reddy et al., 2024) began to incorporate full-length SEC disclosure documents. These works focus on evaluating information retrieval and factual extraction under long-document settings.
Despite incorporating richer contexts, these benchmarks still fundamentally model filings as flattened databases, where core tasks are limited to point-wise fact lookup or single-period detail verification. As a result, they do not test whether models can reason about logical consistency, structural alignment, or semantic evolution across disclosures. To extend beyond single-document retrieval, SEC-QA (Lai et al., 2025) introduces cross-company queries through a template-based generation pipeline. While enabling limited cross-entity comparison, the resulting questions are highly rigid in structure and struggle to capture the heterogeneous and narrative-driven nature of disclosures across different companies. Although SECQUE (BenYoash et al., 2025) further incorporates expert-authored analytical questions spanning tasks such as trend analysis, its comparative settings focus on financial metrics and remain restricted to single reporting periods, leaving non-financial dimensions like risk factors largely unexplored. Beyond task construction, existing benchmarks also suffer from a pronounced black-box issue in their evaluation mechanisms. As illustrated in Table 1, prior work relies solely on answer-level accuracy or coarse-grained labels such as Correct, Partial, or Incorrect. Such shallow evaluation protocols obscure the true failure modes of language models in complex financial reasoning scenarios, making it difficult to distinguish whether errors stem from long-document perception degradation, breakdowns in multi-step financial logic, or misalignment across temporal and structural contexts.
@@ 3.1. Data Source and Statistics
As illustrated in Figure 2, Fin-RATE is built entirely from official filings retrieved via the U.S. SEC EDGAR database [^1] https://www.sec.gov/edgar/search-and-access, ensuring authoritative, traceable, and consistent data sourcing. We select a representative set of filing types that together cover the key dimensions required for realistic financial analysis, including periodic reports (10-K, 10-Q), event-driven disclosures (8-K), governance materials (DEF 14A), foreign issuer updates (6-K), equity incentive plans (S-8), securities offerings (424B series), and shareholder communications (SC series). To support cross-year and cross-company reasoning, we filter companies by: (1) ranking candidates by the number of available 10-K filings and retaining those with sufficient multi-year coverage; and (2) sampling across GICS sectors to ensure broad industry and reporting-style diversity. The collection spans sectors ranging from relatively stable industries such as utilities and consumer goods to dynamic or newly regulated areas such as cryptocurrency services and biopharmaceutical R&D.
All the filings are converted into Markdown, retaining section headings, tables, and lists. We segment each filing with SEC-standard item boundaries into coherent analytical units to support reasoning and retrieval at multiple levels of granularity. The final corpus comprises 15,311 document chunks derived from 2,472 filings, covering 43 companies across 36 industries from 2020 to 2025, with each chunk containing an average of 2,649 words. Periodic and event-driven reports like 10-K, 8-K, and 10-Q forms constitute the majority of the collection, with their distributions and regulatory functions detailed in Appendix A. Based on this corpus, we construct three QA tasks: Detail Reasoning (DR-QA), Enterprise Comparison (EC-QA), and Longitudinal Tracking (LT-QA). Each task contains 2,500 instances. DR-QA questions are grounded in a single chunk, while EC-QA and LT-QA require aggregating evidence from an average of 5.6 and 3.0 chunks, respectively.
@@ 3.2.1. Detail & Reasoning QA (DR-QA)
To ensure the analytical depth of each generated QA pair, we design adaptive prompt templates guided by the official section definitions of various SEC filing types. Prompts for 10-K filings emphasize annual ratio analysis, 10-Q filings highlight quarterly trend dynamics, and 8-K filings direct attention to the quantitative impact of material events. This ensures that each question focuses on the most analytically salient components of the disclosure, effectively reducing the generation of low-value questions and establishing a high-fidelity benchmark for local financial comprehension. See Appendix B for details.
@@ 3.2.2. Enterprise Comparison QA (EC-QA)
To ensure interpretability and regulatory consistency, EC-QA is framed within the same Global Industry Classification Standard (GICS) sector. category and limited to companies filing 10-K reports in the same fiscal year, ensuring consistency in disclosure frameworks and data granularity. During text selection, we pinpoint semantically aligned sections using structured headings from SEC filings, prioritizing high-comparability modules such as “Item 1A: Risk Factors” and “Item 7: Management Discussion and Analysis.” Besides, cross-sectional combinations are introduced to support multi-dimensional comparative questions across business operations, risk exposures, governance structures, and financial performance. To mitigate the risks of information dilution and hallucination caused by long input contexts, EC-QA construction process adopts a two-stage generation strategy. It first generates structured summaries for each selected company section and then constructs comparison-oriented QA pairs from these aligned summaries. In this way, EC-QA balances analytical depth with input complexity control. See Appendix  C for details.
@@ 3.2.3. Longitudinal Tracking QA (LT-QA)
LT-QA is constructed based on long-form disclosures, primarily annual 10-K reports and DEF 14A proxy statements, and adopts a two-stage, summary-first generation pipeline similar to that used in EC-QA. First, structured summaries are generated for each company–year–section unit, capturing key aspects such as strategic orientation. Subsequently, QA instances are constructed based on these summaries, with each question explicitly linking the same section across multiple years and reflecting temporal comparisons and evolutionary reasoning. See Appendix D for details.
@@ 3.3. Rigorous Quality Control
Fin-RATE includes three quality control stages designed to ensure the reliability, factual grounding, and financial relevance of all QAs. In the automated quality control phase, we adopt a dual-model generation-verification mechanism. First, the generation model (DeepSeek-V3.2) produces multiple candidate QAs under either chunk-level or document-level contexts, based on different prompting strategies. Then, two structurally distinct and high-performing LLMs (DeepSeek-V3.2 and Qwen3-235B) independently assess each candidate QA, focusing on whether the answer is sufficiently supported by the context, whether it has numerical or logical errors, and whether it relies on unsupported inference. Based on task-specific validation rules and a standardized rubric, each QA is labeled as Correct, Partial, Incorrect, or Failure. Any QA flagged as erroneous by either model is discarded, and only those rated as correct proceed to the human review stage. In the final manual review phase, each QA that passed automated filtering is confirmed one by one to ensure strict adherence to the original disclosure without subjective interpretation. For each QA type, this process is repeated iteratively until 2,500 valid QAs are retained. Full criteria for manual review are detailed in Appendix E. For QAs involving multiple chunks or cross-document contexts, we examine whether each referenced chunk contributes substantively to the reasoning chain. If any referenced chunk is irrelevant, the QA is rewritten or discarded. For numerical questions, we recalculate all values to avoid potential errors. During this process, we observed recurring error patterns, which motivated the error type taxonomy used in our evaluation.
@@ 4.1.1. Models
  • Finance-Specific Models: To assess the effect of domain adaptation, we evaluate TouchstoneGPT-7B (Wu et al., 2025), Fin-R1 (liu2026finr1largelanguagemodel), Fino1-14B (Qian et al., 2025), and FinanceConnect-13B (CeADAR, 2023).